# Phase transitions at high and low densities for a rotating QCD matter from holography

### v1

@@ 1 Introduction
Holographic approaches provide a natural framework to address this challenge. The AdS/QCD correspondence, derived from the AdS/CFT principles [13, 14, 15], maps strongly coupled, nonperturbative QCD at the boundary to weakly coupled gravity in a five-dimensional AdS5 bulk [16, 17, 18, 19, 20, 21, 22]. Bottom-up models, particularly the soft-wall AdS/QCD model, efficiently incorporate nonperturbative QCD phenomena such as confinement, spontaneous chiral symmetry breaking, and hadronic Regge trajectories [23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. Although QCD lacks exact conformal symmetry, soft-wall constructions keep up with chiral symmetry breaking and the QCD scale with suitable choices of dilaton profiles and warp factors, allowing parameters to match phenomenological data [33, 34, 35, 36, 37]. Finite-density extensions often employ the AdS–Reissner–Nordström approximation [38, 39], but exact solutions such as Andreev’s charged rotating black hole (BH) in AdS5 provide a more accurate holographic description [40]. The fluid/gravity correspondence further connects holographic QCD with dissipative relativistic hydrodynamics describing the QGP [41, 42, 43, 44, 45].
@@
Gauge/gravity duality also reports phase transitions relating the QGP to the hadronic phase [58]. In holographic QCD, the confinement/deconfinement transition is interpreted as a Hawking–Page transition between thermal AdS and an AdS BH [59]. The soft-wall model captures linear Regge trajectories, confinement, and other infrared QCD features [60, 61]. Rotation introduces modified gauge-field dynamics that qualitatively alter BH thermodynamics and the boundary phase structure. In this work, we numerically analyze Hawking–Page transitions in the exact Andreev soft-wall model at finite density and rotation. Using the full on-shell action, including gravitational, dilaton, and Abelian gauge-field contributions, we compute the renormalized action density difference between rotating charged AdS BHs and thermal AdS. This identifies the dominant thermal saddle and maps the QCD-like phase diagram in the $(T,\mu)$ plane for various angular velocities $\omega$. We find three distinct regimes: at zero density, a Herzog-type first-order transition persists with rotation affecting only the critical temperature via relativistic redshift; at high density, the first-order transition survives, but the critical temperature decreases with rotation and the maximal density for Hawking–Page transitions becomes rotation-dependent; at low density and high angular velocity, $\omega l\gtrsim 0.16$, Hawking–Page transitions disappear over a finite chemical potential range, producing smooth crossovers. The interplay of these regimes generates a holographic critical point, whose location is estimated after calibrating the holographic energy scale.
The paper is organized as follows: Sec. 2 contains an overview of the construction of the regularized rotating charged BH action density in AdS space, with its Hawking temperature obtained from the surface gravity formula. In Sections 3, 4, and 5, we describe the first-order transitions that occur for non-rotating matter, at large density and at zero density, respectively. In Sec. 6, we analyzed smooth transitions (crossovers) in the low-density regime, accounting for relativistic rotations. The non-relativistic limit of the phase transitions at low and high densities is studied in Sec. 7, while in Sec. 8 we obtain a numerical estimate for the critical point. Finally, Sec. 9 contains our conclusions.
@@ 2 Rotating charged BH in the exact Andreev’s soft wall model
which shows that the chemical potential transforms as the inverse of the Lorentz factor. From the gauge field in the rotating system (2.12) - (2.14), using the relation (2.16) and assuming $\eta=1$, one obtains the following on-shell version of the $U(1)$ action:
@@ 3 First-order transitions for non-rotating matter
The Hawking temperature is a function of $\mu$, $\omega l$, and $z_{h}$ together. The critical temperatures of deconfinement for rotating matter of a given density can be computed once the critical horizons are known. Using Eqs. (2.6) and (2.16), its dimensionless version in Andreev’s exact soft wall model reads
@@
where the critical horizon appears as a function of the density. In Fig. 1, we have plot $\bigtriangleup\bar{\mathcal{E}}$ as a function $z_{h}$ at $\omega l=0$. As it is shown, the critical horizons ${z_{h}}_{c}$ are sensitive to the chemical potential. Table 1 contains the values of ${z_{h}}_{c}$ at different quark densities, see Appendix A, which were used to compute the corresponding critical temperatures, see Table 1. In Fig. 2, we have plotted $T_{c}$ as a function of $\mu$ for a non-rotating matter according to Table 1. The behavior of $T_{c}$ is similar to that obtained in [74]. The difference here is that we do not apply the Reissner-Nordström (RN) approximation, which consists of taking the limit of small $z$ in the gauge field solution (2.15).
@@
One observes that $T_{c}$ decreases as the quark density increases, until reaching a maximum density value beyond which there are no more transitions and the matter is always described by a plasma. Throughout the curve, the transitions are first-order transitions, in which the Gibbs free energy jumps from one phase to another as the temperature crosses $T_{c}$. The free energies of each phase can be calculated by holographic renormalization; see, for instance, Ref. [66]. For a system in which all particles have zero angular momentum, there is no distinction between the type of transition in the low- and high-density regimes. In non-central heavy-ion collisions, however, the plasma formed exhibits strong vorticity, with angular velocities approaching the speed of light. Such effects must be taken into account in an accurate description of the QCD phase diagram.
@@ 5 First-order transitions at zero density
In the phase diagram, these first-order transitions occur along the temperature axis. This result complies with the transition described by EMD models at zero chemical potential [76]. The $\bigtriangleup\bar{\mathcal{E}}$ curves in Fig. 4 represent this type of transition, with the BH action density crossing the $\bar{z}_{h}$ axis only once at the same point. In the next section, we will carefully analyze the transitions that occur at low (but not zero) densities, between first-order Herzog-type transitions at $\mu=0$ and first-order transitions in the high-density regime that occur at relatively low temperatures, compared to those that occur at low densities, As we will see, there is a non-trivial behavior of the charged BH action densities when subjected to relativistic rotational effects, which are capable of disturbing the stability of QCD matter. Clearly, the phase transitions between the hadronic matter and the QGP in this interval will not be described by first-order transitions for relativistic rotational velocities.
@@ 6 Phase transitions at low densities for relativistic rotational velocities
To visualize this result, we plot the critical temperatures as a function of $\bar{\mu}$ at different rotational velocities in Fig. 8, according to Tables 3, 4, and 5. To the left of the dashed vertical line C, the matter with rotational velocity $\omega l=0.3$ (blue points) is always in the hadronic phase (except for a narrow region near the temperature axis where $\bar{\mu}\approx 0$, see Fig. 5-(H)). This dashed line corresponds to $\bar{\mu}_{EF}(\omega l=0.3)$. Between the dashed lines B and C, there is a region at high temperatures, i. e., for $\bar{T}\geq\bar{T}_{c}(\omega l=0.4)$, where the QGP with rotational velocity $\omega l=0.4$ (yellow points) coexists with the hadronic matter with $\omega l=0.3$. The vertical dashed line B corresponds to $\bar{\mu}_{EF}(\omega l=0.4)$, see Table 3. The same argument is valid for the regions between the vertical dashed lines A, B, and C. This phase-mixing at low energies occurs even for small values of $\bar{\mu}$, as shown in Fig. 7. As $\bar{\mu}$ tends to zero, the mixture will occur due to states that rotate at velocities comparable to the speed of light.
@@
Also from Fig. 7, one concludes that the coexistence between the two phases at high temperatures is limited to a defined region at low densities, until it reaches the maximum value of $\bar{\mu}_{EF}$, i. e., for $\bar{\mu}\approx\bar{\mu}_{EF}(\omega l=0.3)=0.414$. The most critical density is given by the value of $\bar{\mu}$ at $T=0$ for non-rotating matter, which is $\bar{\mu}(T\approx 0)\approx 1.067$. For $\bar{\mu}\geq 1.067$, there are no phase transitions. For $0.414\leq\bar{\mu}\leq 1.067$, the system exhibits first-order transitions. In principle, these transitions also allow the coexistence of the two phases between the $T_{c}$ curves at different rotational velocities, even at very low temperatures, since the values of the maximum critical density at $T=0$ decrease with $\omega l$. This type of coexistence can occur only at temperatures below the $T_{c}(\omega l=0)$ curve, and it is not expected to appear prominently in the QCD phase diagram. At lower temperatures, there is not much energy available for particle kinetics, as particle interactions in highly dense states consume a large portion of the energy. In this case, the transitions will be dominated by first-order transitions for temperatures slightly lower than $T_{c}(\omega l=0)$, and the coexistence between the phases at high densities should occur in an extremely narrow region at very low temperatures, see Fig. 8 – between the blue and yellow $T_{c}$ curves (where the phases are mixed), $T<0.08$, approximately. It decreases even further as rotational velocities increase. We will discuss the non-relativistic limit in the next section.

### v2

@@ 1 Introduction
Holographic approaches provide a natural framework to address this challenge. The AdS/QCD correspondence, derived from the AdS/CFT principles [13, 14, 15], maps strongly coupled, nonperturbative QCD at the boundary to weakly coupled gravity in a five-dimensional AdS5 bulk [16, 17, 18, 19, 20, 21]. Bottom-up models, particularly the soft-wall AdS/QCD model, efficiently incorporate nonperturbative QCD phenomena such as confinement, spontaneous chiral symmetry breaking, and hadronic Regge trajectories [22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. Although QCD lacks exact conformal symmetry, soft-wall constructions keep up with chiral symmetry breaking and the QCD scale with suitable choices of dilaton profiles and warp factors, allowing parameters to match phenomenological data [32, 33, 34, 35, 36]. Finite-density extensions often employ the AdS–Reissner–Nordström approximation [37, 38], but exact solutions such as Andreev’s charged rotating black hole (BH) in AdS5 provide a more accurate holographic description [39]. The fluid/gravity correspondence further connects holographic QCD to dissipative relativistic hydrodynamics that describe the QGP [40, 41, 42, 43, 44].
@@
Gauge/gravity duality also predicts phase transitions that relate the QGP to the hadronic phase [57]. In holographic QCD, the confinement/deconfinement transition is interpreted as a Hawking–Page transition between thermal AdS and an AdS BH [58]. The soft-wall model captures linear Regge trajectories, confinement, and other infrared QCD features [59, 60]. Rotation introduces modified gauge-field dynamics that qualitatively alter BH thermodynamics and the boundary phase structure. In this work, we numerically analyze Hawking–Page transitions in the exact Andreev soft-wall model at finite density and rotation. Using the full on-shell action, including gravitational, dilaton, and Abelian gauge-field contributions, we compute the renormalized action density difference between rotating charged AdS BHs and thermal AdS. This identifies the dominant thermal saddle and maps the QCD-like phase diagram in the $(T,\mu)$ plane for various angular velocities $\omega$. We find three distinct regimes: at zero density, a Herzog-type first-order transition persists with rotation affecting only the critical temperature via relativistic redshift; at high density, the first-order transition survives, but the critical temperature decreases with rotation and the maximal density for Hawking–Page transitions becomes rotation-dependent; at low density and high angular velocity, $\omega l\gtrsim 0.16$, Hawking–Page transitions disappear over a finite chemical potential range, producing smooth crossovers. The interplay of these regimes generates a holographic critical point, whose location is estimated after calibrating the holographic energy scale.
The paper is organized as follows: Sec. 2 contains an overview of the construction of the regularized rotating charged BH action density in AdS space, with its Hawking temperature obtained from the surface gravity formula. In Sections 3, 4, and 5, we describe the first-order transitions that occur for non-rotating matter, at large density and at zero density, respectively. In Sec. 6, we analyze the Gibbs free energy, entropy and specific heat from holographic renormalization, besides smooth transitions (crossovers) in the low-density regime, accounting for relativistic rotations. The non-relativistic limit of the phase transitions at low and high densities is also studied. In Sec. 8 we obtain a numerical estimate for the critical point and pave the comparison with lattice QCD, effective models, and holography in Sec. 8. Finally, Sec. 9 contains our conclusions.
@@ 2 Rotating charged BH in the exact Andreev’s soft wall model
which shows that the chemical potential transforms as the inverse of the Lorentz factor. In particular, we know that $\mu^{\prime}$ corresponds to the chemical potential measured in the comoving (rotating) frame of the plasma and is related to the boundary chemical potential $\mu$ through a Lorentz factor. This relation ensures consistency with relativistic thermodynamics. Although our phase diagram is expressed in terms of $\mu$, the introduction of $\mu^{\prime}$ provides a useful interpretation of density effects in rotating systems. From the gauge field in the rotating system (2.12) - (2.14), using the relation (2.16) and assuming $\eta=1$, one obtains the following on-shell version of the $U(1)$ action:
@@ 3 First-order transitions for non-rotating matter
The Hawking temperature is a function of $\mu$, $\omega l$, and $z_{h}$ together. The critical temperatures for deconfinement of rotating matter at a given density can be computed once the critical horizons are known. Using Eqs. (2.6) and (2.16), its dimensionless version in Andreev’s exact soft wall model reads
@@
where the critical horizon appears as a function of the density. In Fig. 1, we have plot $\bigtriangleup\bar{\mathcal{E}}$ as a function $z_{h}$ at $\omega l=0$. As it is shown, the critical horizons ${z_{h}}_{c}$ are sensitive to the chemical potential. Table 1 lists the values of ${z_{h}}_{c}$ at different quark densities (see Appendix A), which were used to compute the corresponding critical temperatures (see Table 1). In Fig. 2, we have plotted $T_{c}$ as a function of $\mu$ for a non-rotating matter according to Table 1. The behavior of $T_{c}$ is similar to that obtained in [73]. The difference here is that we do not apply the Reissner-Nordström (RN) approximation, which consists of taking the limit of small $z$ in the gauge field solution (2.15).
@@
One observes that $T_{c}$ decreases with increasing quark density, until it reaches a maximum density beyond which no further transitions occur and the matter is always described by a plasma. Throughout the curve, the transitions are first-order transitions, in which the Gibbs free energy jumps from one phase to another as the temperature crosses $T_{c}$. The free energies of each phase can be calculated by holographic renormalization; see, for instance, Ref. [65]. For a system in which all particles have zero angular momentum, there is no distinction between the type of transition in the low- and high-density regimes. In non-central heavy-ion collisions, however, the plasma formed exhibits strong vorticity, with angular velocities approaching the speed of light. Such effects must be accounted for in an accurate description of the QCD phase diagram.
@@ 5 First-order transitions at zero density
In the phase diagram, these first-order transitions occur along the temperature axis. This result is consistent with the transition predicted by EMD models at zero chemical potential [75]. The $\bigtriangleup\bar{\mathcal{E}}$ curves in Fig. 4 represent this type of transition, with the BH action density crossing the $\bar{z}_{h}$ axis only once at the same point. In the next section, we will carefully analyze the transitions that occur at low (but not zero) densities, between first-order Herzog-type transitions at $\mu=0$ and first-order transitions in the high-density regime that occur at relatively low temperatures, compared to those that occur at low densities, As we will see, there is a non-trivial behavior of the charged BH action densities when subjected to relativistic rotational effects, which are capable of disturbing the stability of QCD matter. Clearly, the phase transitions between the hadronic matter and the QGP in this interval will not be described by first-order transitions for relativistic rotational velocities.
@@ 6 Phase transitions at low densities for relativistic rotational velocities
To visualize this result, we plot the critical temperatures as a function of $\bar{\mu}$ at different rotational velocities in Fig. 8, according to Tables 3, 4, and 5. To the left of the dashed vertical line C, the matter with rotational velocity $\omega l=0.3$ (blue points) is always in the hadronic phase (except for a narrow region near the temperature axis where $\bar{\mu}\approx 0$, see Fig. 5-(H)). This dashed line corresponds to $\bar{\mu}_{EF}(\omega l=0.3)$. Between the dashed lines B and C, there is a region at high temperatures, i. e., for $\bar{T}\geq\bar{T}_{c}(\omega l=0.4)$, where the QGP with rotational velocity $\omega l=0.4$ (yellow points) coexists with the hadronic matter with $\omega l=0.3$. The vertical dashed line B corresponds to $\bar{\mu}_{EF}(\omega l=0.4)$, see Table 3. The same argument applies to the regions between the vertical dashed lines A, B, and C. This phase-mixing at low energies occurs even for small values of $\bar{\mu}$, as shown in Fig. 7. As $\bar{\mu}$ tends to zero, the mixture will occur due to states that rotate at velocities comparable to the speed of light.
@@
Also from Fig. 7, one concludes that the coexistence between the two phases at high temperatures is limited to a defined region at low densities, until it reaches the maximum value of $\bar{\mu}_{EF}$, i. e., for $\bar{\mu}\approx\bar{\mu}_{EF}(\omega l=0.3)=0.414$. The most critical density is given by the value of $\bar{\mu}$ at $T=0$ for non-rotating matter, which is $\bar{\mu}(T\approx 0)\approx 1.067$. For $\bar{\mu}\geq 1.067$, there are no phase transitions. For $0.414\leq\bar{\mu}\leq 1.067$, the system exhibits first-order transitions. In principle, these transitions also allow the coexistence of the two phases between the $T_{c}$ curves at different rotational velocities, even at very low temperatures, since the values of the maximum critical density at $T=0$ decrease with $\omega l$. This type of coexistence can occur only at temperatures below the $T_{c}(\omega l=0)$ curve, and it is not expected to appear prominently in the QCD phase diagram. At lower temperatures, there is little energy available for particle kinetics, as particle interactions in highly dense states consume a large fraction of the energy. In this case, the transitions will be dominated by first-order transitions for temperatures slightly lower than $T_{c}(\omega l=0)$, and the coexistence between the phases at high densities should occur in an extremely narrow region at very low temperatures, see Fig. 8 – between the blue and yellow $T_{c}$ curves (where the phases are mixed), $T<0.08$, approximately. It decreases even further as rotational velocities increase. We will discuss the non-relativistic limit in the next section.